OpenStudy (anonymous):

How does ln(5) - ln(4) - ln(6) + ln(5) simplify to ln(25/24). I understand that ln5-ln4 = ln(5/4), not sure if that helps to simplify, but I don't see it.

OpenStudy (butterflydreamer):

do you know your log laws? :) If you SUBTRACT ln(x) - ln (y) = ln (x/y) because you DIVIDE. If you ADD ln (x) + ln (y) = ln (xy) because you MULTIPLY. So theeeenn ln (5) - ln (4) - ln (6) + ln (5) = ln ( 5 divided by 4 divided by 6 times 5) So to break it down: ln (5) - ln (4) = ln (5/4) right? Then ln (5/4) - ln (6) = ln (5/24) Then ln (5/24) -ln (5) = ln (5/24 * 5) = ln (25/24)

OpenStudy (anonymous):

ah very nice :), ty @butterflydreamer, and all else

OpenStudy (butterflydreamer):

you're welcome! :) Hope you get it xD Looks kinda confusing

OpenStudy (legends):